OpenStudy (anonymous):

Ok, this might be a stupid question, but how do you calculate the density of CH4 at .98 atm and 35 degrees Celsius? I know that Density is Mass/Volume but in this case how would you find the Mass and Volume?

OpenStudy (anonymous):

You don't. Assuming ideal gas, \[PV=nRT \\ \\ \frac{P}{RT}=\frac{n}{V} \\ \\ \frac{P}{RT}=\frac{m}{M_rV} \\ \\ \frac{PM_r}{RT}=\rho\]

OpenStudy (anonymous):

How did you get from PV = nRT to \[\frac{ P }{ RT }= \frac{ n }{ V }\] did you first divide both sides by RT \[\frac{ PV }{ RT }= n\] Then divide both sides by V \[\frac{ P }{ RT }=\frac{ n }{ V }\] I don't follow though how you replaced n with m and what is \[M_{r}\]